OpenStudy (anonymous):

Please help! The line passes through (2,-4) and is perpendicular with the slope of -2

OpenStudy (anonymous):

Given the slope is -2 or -2/1 do you know how to find the perpendicular slope?

OpenStudy (anonymous):

I don't know how to find the perpendicular slope - Is it the point slope formula? and thank you! =)

OpenStudy (anonymous):

turn -2/1 upside down and get rid of the negative sign to give us 1/2

OpenStudy (anonymous):

is that a negative reciprocal and why does it have to be a negative reciprocal?

OpenStudy (anonymous):

yes it is a negative reciprocal and it needs to be a negative reciprocal because it is perpendicular

OpenStudy (anonymous):

OHHHH - the answer key says the answer is y=1/2 x - 5 -- How did they get that answer - especially the -5 part?? Thank you so much!

OpenStudy (anonymous):

Are you okay with the 1/2 part??

OpenStudy (anonymous):

yes - got that part. I just don't understand the y intercept part (-5). :)

OpenStudy (anonymous):

We can use y = mx + b were y = -4 x = 2 m = 1/2 Substitute -4 = (1/2)(2) + b then multiply -4 = 1 + b subtract 1 from both sides -5 = b

OpenStudy (anonymous):

i got it.....you have to use the point slope formula with the negative reciprocal slope!! right? :)

OpenStudy (anonymous):

You can, but I did not ...

OpenStudy (anonymous):

o,ok...well you're a smarty :) thank you so much for your help!